OpenStudy (anonymous):

Solve for x. 7(x - 3) + 3(4 - x) = -8 x = 1/4 x = 25/10 x = -16/19 x = -41/10

OpenStudy (anonymous):

1/4, need the work?

OpenStudy (anonymous):

no thank you

Parth (parthkohli):

7x - 21 + 12 - 3x = 8 => 4x - 9 = -8 => 4x = 1 => x = 1/4

OpenStudy (anonymous):

its like the same, u did b4..